# Searching for internal symbols underlying deep learning

### v1

@@ 3.2 How confident are you, DNNs?
In the analysis above, we used $CM(i,j)$ to infer the most probable class by using symbol as a hash key and $CM(i,j)$ as a look-up table. However, this look-up table can be read in opposite direction. If we know the label $j$ of inputs, we can infer how likely individual symbols $S_{i}$ appear in DNNs. With this idea in mind, we defined ‘Expected Symbol Score (ESS)’ per layer with respective label/class $j$ (Eq. 5).
@@
, where $S_{i}$ denotes the indices of 9 symbols obtained from each ROI. That is, $ESS_{j}$ with respective class $j$ is the mean value of 9 probability from $P(i,j)^{m}$; see Eq. (5). We estimated $ESS_{j}$ from all 4 layers. First, we asked if ESSs obtained from layer 4 can be correlated with the accuracy of DNNs’ answers. Specifically, we split ESSs into two distributions depending on whether or not DNNs made the correct decisions.
We found that ESSs were significantly higher when predictions were correct (see data marked by * in Fig. 4A) than when they were incorrect (see data without * in Fig. 4A). To quantify the separation between correct and incorrect predictions, we estimated AUROC (the orange line in Fig. 4B), suggesting that ESSs can indicate DNNs’ likelihood of making correct predictions. We also examined ESSs from layers 1, 2 and 3 and AUROC between correct and incorrect answers (Fig. 4B). Further, we examined the norm of ESSs $||(ESS_{1},ESS_{2},ESS_{3},ESS_{4})||$ from all layers, and AUROC estimated from this norm were similar to those obtained from layer 4 (see blue line Fig. 4B).
Figure 4: Symbol-based confidence level. (A), Expected symbol score (ESS) in ResNet18(Res18), ResNet50(Res50), VGG19 and DenseNet121(DENSE). The asterisks denote ESSs when DNNs make correct predictions, which are higher than those when DNNs make incorrect predictions. (B), AUROC between ESS when DNNs are correct and ESS when DNNs are wrong. We tested ESS from layer 4(orange), layer 3(green), layer2(red) and layer1(violet). Also we tested the norm from all layers, and AUROC between the norms is shown in blue.
@@ 3.3 Detecting out-of-distribution (OOD) examples
We assumed that hidden layer responses (i.e., symbols), highly optimized for in-distribution examples during training, are in synchronization across layers during inference, when inputs are drawn from in-distribution, but they can be out of synchronization with each other, when inputs are drawn from out-of-distribution. Further, we speculated that ESSs can be used to measure the synchrony between layers and thus detect OOD examples. To test this line of hypotheses, we extracted the symbols from NINCO dataset [33], which contain OOD examples carefully curated against ImageNet.
@@
Fig. 5A visualizes ESSs ($ESS_{1},ESS_{2},ESS_{3}$) from layers 1, 2, 3 in a 3D space. Blue circles and orange triangles denote in-distribution and OOD examples respectively. As shown in the figure, ESSs were significantly different between the in-distribution and OOD examples, and we quantified this observation by estimating AUROC. The estimated AUROC from ESSs of all 3 layers and their norms $||(ESS_{1},ESS_{2},ESS_{3})||$ suggest that ESSs are well separated between normal and OOD examples and can be used to detect OOD examples.
@@
Figure 5: (A), Comparison of ESSs between in-distribution and OOD examples. Orange triangles denote ESSs from layers 1,2 and 3 when inputs are drawn from OOD. Blue circles denote ESSs when inputs are drawn from in-distribution examples. (B), AUROC between ESSs for all 4 DNNs. AUROC were calculated using ESSs from layers 3, 2 and 1. Additionally, ESSs were also calculated using the norm of ESS from all 3 layers.
3.4 Detecting adversarial examples
@@
Since the discovery of DL models’ vulnerability to adversarial examples [34, 35], adversarial perturbation has emerged as one of the greatest threats to DL’s safety. Given that all inputs are progressively processed in modern DNNs, it seems natural to assume that adversarial perturbation may change internal representation on layer-by-layer basis. An earlier study [36] found that the difference in representations in early layers was not noticeable, but in late layers it grew significantly. We asked if ESSs can be used to detect the influence on adversarial perturbation. Specifically, we crafted adversarial examples from validation images of Mixed_13 using AutoAttack [37], which is known to be one of the most effective attacks. In the experiment, we used a ‘standard’ mode with $eps=0.03$ and $norm=Linf$. For individual inputs, we used labels of clean images to detect ROIs and estimated the corresponding symbols; that is, ROIs are identified with original labels, not manipulated predictions.
For ROIs, we estimated the norm of ESSs from all 4 layers $||(ESS_{1},ESS_{2},ESS_{3},ESS_{4})||$ and calculated AUROC by comparing the norms of ESSs between normal and adversarial inputs. We note that the AUROC estimated from the norms were close to 0.5 (see the blue line in Fig. 6), suggesting that the symbols in DNNs were not significantly different between normal and adversarial examples. Consistent with this finding, the symbols of adversarial inputs in layer 4 predict the original classes as accurately as the symbols of the normal inputs do (the thick green line in Fig. 6). These results suggest that the symbols are much more resilient than the models’ final predictions. That is, the detection of adversarial inputs may require methods other than comparing internal symbols.
@@
So far, ESSs have been computed using the most probable class predicted by the layer 4 correlation map $CM(i,j)$, but they can also be computed using DNNs’ predictions. If ESSs are computed with DNNs’ predictions, we can effectively evaluate the consistency between models’ decisions and internal symbols, which we assume could be useful in detecting adversarial perturbation. However, we analyzed unmodified ImageNet models, whose output nodes correspond to 1,000 classes, not 78 of Mixed_13, Consequently, the predictions were not necessarily confined to 78 classes of Mixed_13. Thus, to build ESSs using their predictions, we selected adversarial and normal symbols when predictions are one of 78 classes (see Table 3 for the number of ROIs included in this analysis). This Decision-based ESSs are compared in Fig. 6 (see the orange line). As shown in the figure, Decision-based ESSs are significantly different for all 4 models, suggesting that ESSs with respect to DNNs’ (manipulated) predictions can be used to detect adversarial perturbation.
@@
Figure 6: Symbol-based detection of adversarial perturbations. The orange line denotes AUROC calculated from decision-based ESS, whereas the blue dashed line denotes AUROC calculated ESS. The green line denotes the accuracy of symbol-based predictions on the adversarial inputs.
@@ 3.5 Temporary learning
One of DL’s shortcomings is the lack of continual learning ability, whereas the brain can learn continuously. The brain can use previous knowledge to incorporate new knowledge, but DL models’ catastrophic forgetting, which deletes previously obtained information to learn the new, prevents continual learning. Since symbols can serve as previous knowledge, we speculate that the temporary mapping between symbols of in-distribution examples and OOD examples may allow a short-term (temporary) learning of OOD examples, which can serve as a stop-gap measure, until a more formal retraining becomes available. To address this possibility, we randomly split NINCO symbols (obtained from 18 classes) into 2 distinct distributions, which are training and test sets. The size of the training and test sets are half of the full size of NINCO symbols. Then, the training symbols were used to construct $CM(i,j)$, and we asked if the test symbols can predict the inputs’ classes. As selecting training/test symbols is done randomly, we resampled them 100 times and reported the results from all 100 experiments (Fig. 7). We observed notable differences between the models but do not know the reasons behind them. More importantly, the accuracy is around 40-50% (Fig. 7) for all 4 models, supporting that internal symbols can be used for short-term learning.
Figure 7: The accuracy of predictions on the OOD examples for ResNet18(Res18), ResNet50(Res50), VGG19 and DenseNet121(DENSE). We resampled trainig and test examples from 18 classes of NINCO dataset [33] for 100 times, and collected their accuracies.
@@ 4 Discussion
We probed the existence of internal symbols closely linked to DNNs’ decision making. More specifically, we combined STCert and traditional machine learning algorithms to detect internal symbols. Our analysis of these internal symbols suggest that they can be used to 1) predict levels of accuracies of DNNs’ answers, 2) detect OOD and adversarial inputs and 3) enable temporary learning of OOD examples. Based on our analysis, we propose that decoding of internal symbols could play a crucial role in making DNNs’ operations more reliable and safer.
@@ 4.1 Limitations
We should underline that the purpose of our study is not to provide a comprehensive survey of methods that can extract internal symbols and potential utilities of internal symbols. Instead, we focused on searching for positive evidence of the existence of internal symbols associated with DNNs’ decisions and exploring their applications. Below, we list the limitations of our analysis as potential guides for future studies on internal symbols.
  • We selected Mixed_13 and ImageNet models for two reasons. First, if we extract symbols from poorly trained models, the analysis can be biased. Second, as the training set of ImageNet is huge, analyzing them requires substantial computing resources. As Mixed_13 contains images from well-conceived 13 visual categories and balanced number of classes across all categories, it is a good subset to analyze the internal symbols related to ImageNet. However, this raises the question of scalability of our analysis, and we will address this in the future.
  • We extracted symbols on layer-by-layer basis and assumed that they are encoded collectively by neurons with the same receptive fields across channels. However, the feature visualization shows that some single neurons can encode complex features recognizable to humans. We speculate that some symbols are encoded by a set of neurons, and others, by single neurons, but this needs further study.
  • Our analysis searched for symbols related to labels of ImageNet images, but visual objects rarely exist alone in visual scenes. Other objects or background can serve as context to an object of interest. That is, the symbols can interact with one another. Such functional links between symbols are ignored in our analysis.
  • We examined convolutional networks (CNNs) pretrained on ImageNets. Although CNNs are one of the popular architectures, transformers have been widely used, especially in language models. As transformers are based on fully connected layers and self-attention modules, it remains unclear how to probe internal symbols relevant to their decision-making.
4.2 Broader Impacts
We propose a potential algorithm to analyze DL models’ decision. As this algorithm aims to reduce DL models’ mistakes, we expect it could be helpful in making DL more reliable and are not aware of any potential negative impacts.
Appendix A Appendix and Supplementary Material
A.1 STCert implementation
The two versions of STCert were discussed in the original study [20]. The naive STCert uses the bounding boxes returned by the segmentation models as ROIs. The context-aware STCert adds some backgrounds to the bounding boxes to create ROIs. As the context-aware STCert is more robust, we adopted the context-aware STCert in this study. The workflow of STCert can be summarized as follows:
  1. 1.
    Detect bounding boxes.
  2. 2.
    Create 5 candidates of bounding boxes by enlarging ROIs. The first candidate is ROI itself.
  3. 3.
    Crop pixels within the candidates and forward them into DNNs. That is, we have the list of 5 second thought predictions.
  4. 4.
    Examine if the original prediction is included in the list of second predictions.
  5. 5.
    Return the bounding box as ROI, if the original prediction is in the list. If not, ignore the bounding box and return Null.
A.2 Potential links to the brain structure
Our analysis raises the possibility that internal symbols can be used to augment DNNs’ decision-making process. It may even be possible for a second network to use symbols to correct responses in each layer locally during inference. When immediate decisions are necessary, DNNs’ decisions can be used, as they are, but for more reliable answers, a second network may use symbols to modify hidden layers’ responses. This may remind some readers of thalamo-cortical loops of the brain. Our analysis has been inspired by thalamo-cortical loops and indicates that thalamo-cortical loop-like error correction could be constructed using internal symbols to build more reliable DNNs.
@@ A.2 Potential links to the brain structure
Supplementary Figure 1: The correlations between the symbols and classes. Top, middle and bottom rows show the correlations observed from ResNet50, VGG19 and DenseNet121. $y$-axis denotes the indices of symbols, $x$-axis denotes the class. (A)-(D), the correlations observed in layers 1-4, respectively.
Supplementary Figure 2: The correlations between symbols and classes in ResNet18. These plots show the correlation maps between 200 consecutive symbols and the label of inputs observed in layer 1 (A), layer 2 (B), layer 3 (C) and layer 4 (D). We note that some consecutive symbols are correlated with the same classes and thus can be merged together, especially in layer 4.

### v2

@@ 3.2 How confident are you, DNNs?
In the analysis above, we used $CM(i,j)$ to infer the most probable class by using the symbol indice as a hash key and $CM(i,j)$ as a look-up table. However, this look-up table can be interpreted in a completely opposite direction. If we know the label $j$ of inputs, we could infer the likelihood of individual symbols $S_{i}$ appearing in DNNs. With this in mind, we defined ‘Expected Symbol Score (ESS)’ per layer with respective label/class $j$ (Eq. 5).
@@
, where $S_{i}$ denotes the indices of 9 symbols obtained from each ROI. That is, $ESS_{j}$ with respective class $j$ is the mean value of 9 probabilities from $P(i,j)^{m}$; see Eq. (5). If all observed symbols $S_{i}$s are exclusively correlated with class $j$, $P(i,j)=1$, $P(i,k\neq i)=0$, we can expect that $ESS_{j}=1$, and $ESS_{k\neq i}=0$. By contrast, if symbols contain no information related to classes, we can expect that $P(i,j)=ESS_{j}=1/78$, where 78 denotes the number of classes.
With this possibility in mind, we asked if ESSs obtained from layer 4 could be correlated to the accuracy of DNNs’ answers. Specifically, we split ESSs into two distributions depending on DNNs’ ability to produce correct decisions. We found that ESSs were significantly higher when the predictions were correct (see data marked by * in Fig. 6A) than when they were incorrect (see data without * in Fig. 6A). To quantify the separation between correct and incorrect predictions, we estimated AUROC (the orange line in Fig. 6B), suggesting that ESSs indicate DNNs’ likelihood of making correct predictions. We also examined ESSs from layers 1, 2 and 3 and AUROC between correct and incorrect answers (Fig. 6B). Further, we examined the norm of ESSs $||(ESS_{1},ESS_{2},ESS_{3},ESS_{4})||$ from all layers, and AUROC estimated from this norm were similar to those obtained from layer 4 (see blue line Fig. 6B).
Figure 6: Symbol-based confidence level. (A), Expected symbol score (ESS) in ResNet18(Res18), ResNet50(Res50), VGG19, DenseNet121(DENSE). ViT and ResNet50 trained on Oxford IIT-PET (PET). The asterisks denote ESSs when DNNs make correct predictions, which are higher than those when DNNs make incorrect predictions. (B), AUROC between ESS when DNNs are correct, and ESS when DNNs are wrong. We tested ESS from layer 4 (orange), layer 3 (green), layer2 (red) and layer1 (violet). Also we tested the norm from all layers, and AUROC between the norms is shown in blue.
@@ 3.3 Detecting out-of-distribution (OOD) examples
We assumed that hidden layer responses (i.e., symbols), highly optimized for in-distribution examples during training, are in synchronization across layers during inference, when inputs are drawn from in-distribution but that they are out of synchronization when inputs are drawn from out-of-distribution. Further, we speculated that ESSs could be used to measure the level of synchrony between layers and thus detect OOD examples. To test this line of hypotheses, we extracted the symbols from NINCO dataset [35], which contain OOD examples carefully curated against ImageNet.
@@
Fig. 7A visualizes ESSs ($ESS_{1},ESS_{2},ESS_{3}$) from layers 1, 2, 3 of ResNet18 in a 3D space, in which $x,y,z$ denote $ESS_{1},ESS_{2},EES_{3}$, respectively. The orange circles and the blue triangles denote in-distribution and OOD examples respectively. As shown in the figure, ESSs were substantially different between in-distribution and OOD examples, and we quantified this observation by estimating AUROC for all 6 models. The estimated AUROC from ESSs of all 3 layers and their norms $||(ESS_{1},ESS_{2},ESS_{3})||$ suggest that ESSs are effectively separated between normal and OOD examples and can be used to detect OOD examples.
@@
Figure 7: (A), Comparing ESSs between in-distribution and OOD examples. The orange triangles denote ESSs from layers 1,2 and 3 when inputs are drawn from OOD. The blue circles denote ESSs when inputs are drawn from in-distribution examples. (B), AUROC between ESSs for all 6 moels. AUROC were calculated using ESSs from layers 3, 2 and 1. Additionally, ESSs were also calculated using the norm of ESS from all 3 layers.
3.4 Symbols used to address the vulnerabilities to adversarial perturbations
@@
Since the discovery of DL models’ vulnerability to adversarial examples [36, 37], adversarial perturbation has emerged as one of the greatest threats to DL safety. Given that all inputs are progressively processed in modern DNNs, it seems natural to assume that adversarial perturbation may change internal representation on layer-by-layer basis. An earlier study [38] found that the difference in representations in the early layers was not noticeable, but in the late layers it grew significantly. We asked if ESSs can be used to detect the influence on adversarial perturbation. Specifically, we crafted adversarial examples from validation images of Mixed_13 using AutoAttack [39], which is known to be one of the most effective attacks. In the experiment, we used a ‘standard’ mode with $eps=0.03$ and $norm=Linf$. For individual inputs, we used labels of clean images to detect ROIs and estimated the corresponding symbols; that is, ROIs are identified with original labels, not the manipulated predictions.
For ROIs, we estimated the norm of ESSs from all 4 layers $||(ESS_{1},ESS_{2},ESS_{3},ESS_{4})||$ and calculated AUROC by comparing the norms of ESSs between normal and adversarial inputs. We note that the AUROC estimated from the norms were close to 0.5 (see the orange bars in Fig. 8), suggesting that the symbols in DNNs were not significantly different between normal and adversarial examples. Thus, we tested an alternative method. So far, ESSs have been computed using the most probable class predicted by the layer 4, which is the closest to the logit/output layer in the model. Specifically, we used layer 4’ s symbols ($S_{i}$) and correlation map $CM(i,j)$ to predict a class label $j$, but ESSs can also be computed using DNNs’ predictions directly. That is, we can set $j$ to be the model’s prediction.
@@
If ESSs are computed with DNNs’ predictions, we can effectively evaluate the consistency between models’ decisions and internal symbols, which we assume could be useful in detecting adversarial perturbation. Notably, we analyzed unmodified ImageNet models, whose output nodes correspond to 1,000 classes, not 78 of Mixed_13. As we did not set the target class during adversarial attacks, the predictions were not necessarily confined to 78 classes of Mixed_13. Thus, to build ESSs using their predictions, we selected adversarial and normal symbols when predictions are one of 78 classes in Mixed_13 (see Table 4 for the number of ROIs included in this analysis). For ResNet50 trained on Oxford-IIT PET dataset (shown as ‘PET’ in the $x$-axis), we used all perturbed inputs. Decision-based ESSs are compared between normal and adversarial inputs (see the blue bars in Fig. 8). As shown in the figure, ‘decision-based ESSs’ are significantly different between normal and adversarial inputs in all 6 models, suggesting that ESSs with respect to DNNs’ (manipulated) predictions can be used to detect adversarial perturbation.
@@
Figure 8: Symbol-based detection of adversarial perturbations. We estimated ESS using models’ decisions and the predictions of final layer (i.e., L4). Then, we tested if ESS are significantly different between normal and adversarial inputs. The orange line denotes AUROC calculated from decision-based ESS, whereas the blue dashed line denotes AUROC calculated ESS. The green line denotes the accuracy of symbol-based predictions on the adversarial inputs.
Further, inspired by the results suggesting that symbols can be used to predict labels of inputs (Fig. 5), we made predictions on adversarial inputs using symbols in L4 (e.g., $12^{th}$ embedding layer for ViT). Fig. 9, compares the accuracy of symbol-based predictions and the accuracy of the models’ original predictions on adversarial inputs. We note that even though adversarial inputs effectively lowered the accuracy of DL models’ predictions (see orange bars), their impact on the symbol-based predictions was minimal. This result indicates that symbols can be used to address DL models’ vulnerabilities to adversarial perturbations.
Figure 9: The accuracy of symbol-based predictions on adversarial inputs perturbations. We first made predictions using the symbols in final layer (i.e., L4) and evaluated their accuracy. Blue bars show that the symbols in L4 can accurately predict the original classes of adversarial inputs. For comparison, we also display the accuracy of DL models’ original predictions (see orange bars).
@@ 3.5 Temporary learning
One of DL’s shortcomings is the lack of continual learning ability, whereas the brain can learn continuously. The brain can use previously obtained knowledge and incorporate new knowledge, but DL models’ catastrophic forgetting, the deletion of previously obtained information to learn the new, prevents continual learning. Since symbols can serve as previous knowledge, we speculate that the temporary mapping between symbols of in-distribution examples and OOD examples may allow a short-term (temporary) learning of OOD examples, which can serve as a stop-gap measure, until a more formal retraining becomes available. To address this possibility, we randomly split NINCO symbols (obtained from 18 classes) into 2 distinct distributions, training sets and test sets. The size of the training and test sets are half the size of NINCO symbols. Then, the training symbols were used to construct $CM(i,j)$, where class label $j$ is one of 18 classes of NINCO dataset. Then, we asked if the test symbols can predict the inputs’ classes. As training/test symbols are selected randomly, we resampled them 100 times and reported the results from all 100 experiments (Fig. 10). We observed notable differences between the models. More importantly, the accuracy is around 40-50% (Fig. 10) for all 6 models, which is far higher than the random guess ($1/18$), supporting that internal symbols can be used for short-term learning.
Figure 10: The accuracy of predictions on the OOD examples for ResNet18(Res18), ResNet50(Res50), VGG19, DenseNet121(DENSE) and ViT. We resampled the training and test examples from 18 classes of NINCO dataset [35] for 100 times and collected their accuracies.
3.6 Symbols associated with internal features
In this study, we extract symbols from ROIs. This is based on the assumption that classifiers rely on internal features of visual objects. We tested this assumption by creating cropped versions of ROIs (containing objects) and evaluated DL models’ accuracy on full images, ROIs and cropped ROIs. Specifically, the three cropped versions, ROI81, ROI64 and ROI49%, contain 81%, 64% and 49% of pixels of ROIs obtained from STCert. All cropped ROIs and the full ROI are aligned to have the same centers, and we tested DL models’ accuracy on the test set of Mixed_13. The accuracy on ROI is normalized to the accuracy on full images. We note that the drop in the accuracy on ROIs is limited even at ROI:49% (Table 5), suggesting that DL models can indeed learn to use internal features of visual objects, which supports our assumption.
Table 5: We tested the accuracy of 6 ImageNet models, ResNet18(RES18), ResNet50(RES50), VGG19, DenseNet121(DENS121) and VIT.
Model Full ROI ROI81 ROI64 ROI49
RES18 1 0.851 0.835 0.819 0.79
RES50 1 0.896 0.88 0.87 0.852
VGG19 1 0.839 0.817 0.799 0.773
DEN121 1 0.87 0.848 0.829 0.813
VIT 1 0.896 0.879 0.868 0.851
@@ 4 Discussion
We probed the existence of internal symbols closely linked to DNNs’ decision making. More specifically, we combined STCert and traditional machine learning algorithms to detect internal symbols. Our analysis of these internal symbols suggest that they can be used to 1) predict levels of accuracies of DNNs’ answers (Fig.5), 2) detect OODs (Fig.7) and adversarial inputs (Fig.8), 3) make robust predictions on adversarial inputs (Fig. 9) and 4) enable temporary learning of OOD examples (Fig.10). Based on our analysis, we propose that internal symbols could play a crucial role in addressing the shortcomings of DNNs’ and helping us build more reliable and safer DL models.
@@ 4.1 Limitations
We should underline that the purpose of our study is not to provide a comprehensive survey of methods that can extract internal symbols and potential utilities of internal symbols. Instead, we focused on searching for positive evidence of the existence of internal symbols associated with DNNs’ decisions and exploring their applications.